OpenStudy (snowcrystal):

:( can someone help me....I don't understand this question... Please help The given lengths are two sides of a right triangle. All three side lengths of the triangle are integers, and together they form a Pythagorean triple. Find the length of the third side, then indicate whether it is a leg or a hypotenuse. 30 and 40 15 and 36 70 and 250 45 and 51 15 and 20 96 and 100

OpenStudy (campbell_st):

so you are being asked if these 2 sides of a right triangle are part of a triangle that has a hypotenuse that is a whole number as an example.. pythagoras' theorem is basically \[a^2 + b^2 = c^2\] so a right triangle with shorter sides 5 and 12... will the hypotenuse be an integer of whole number then \[c^2 = 5^2 + 12^2~~~so~~~c^2 = 25 + 144 ~~~~~c^2 = 169\] now take the square root of both sides to find c, the hypotenuse \[c = \sqrt{169} ~~~~~~c = 13\] so right triangle with sides 5, 12 and 13 meets the rule set in the question. hope it helps

OpenStudy (snowcrystal):

That helps a lot. How do i figure out if it is a leg or hypotenuse though?

OpenStudy (snowcrystal):

@campbell_st

OpenStudy (campbell_st):

well you may need 2 calculations.... I looked at the question and read it as you were given the shorter legs... the alternative is to use subtraction \[b^2 = c^2 - a^2\] c is the larger of the pair...a is the smaller the 1st and 2nd last will need to adding version